FREQUENCY COMB-REFERENCED MEASUREMENTS OF SELF- AND NITROGEN-PERTURBED LINE SHAPE PARAMETERS IN THE ν_{1} + ν_{3} BAND OF ACETYLENE

\underline{MATTHEW J. CICH}, GARY V. LOPEZ, TREVOR J. SEARS^{a}, Department of Chemistry, Stony Brook University, Stony Brook, New York 11794; C. P. MCRAVEN, Department of Chemistry, Brookhaven National Laboratory, Upton, New York 11973; A. W. MANTZ, Department of Physics, Astronomy, and Astrophysics, Connecticut College, New London, CT 06320; and DANIEL HURT-MANS, Service de Chimie Quantique et de Photophysique(Atoms, Molecules et Atmospheres), Universite Libre de Bruxelles, Bruxelles, Belgium B-10050.

Using an extended cavity diode laser locked to a single component of an Er-fiber-based femtosecond frequency comb, we have made precise measurements of absorption spectral line shapes in a temperature controlled cell. Varying pressures of acetylene and nitrogen were used to determine the N_{2} pressure-dependent parameters for the P(11) line in the ν_{1} + ν_{3} combination band of acetylene at 195 739.649 513(8) GHz. The temperature dependence of the line shape was determined from measurements at several temperatures, varying from 296 K to 125 K. With the absolute frequency positions at each point on the frequency scale determined by the comb, each experimental data set has better than 10^{−4} fractional error. Parameters describing the line shape, such as pressure-dependent broadening, narrowing and shift coefficients, can be obtained with standard deviations less that 0.1%. The data have been used to test various line shape models beyond the standard Voigt approximation including those with narrowing parameters (Rautian and Galatry models) and those with speed-dependence (Speed-dependent Voigt and Speed-dependent Nelkin-Ghatak models). Fitting results will be presented and the relative performance of the models will be discussed. Acknowledgements:Acknowledgement is made to the Donors of the American Chemical Society Petroleum Research Fund for partial support of this research. CPM gratefully acknowledges support by DOE EPSCoR grant DOE-07ER46361 for work conducted at the University of Oklahoma. The measurements and analyses were performed under grants NNX09AJ93G and NNX08AO78G from the NASA Planetary and Atmospheres program.

aalso: Department of Chemistry, Brookhaven National Laboratory, Upton, New York 11973